### v3

@@ 4.1 AC4 Framework
For quadratic circuits, we solve the coefficients and verify if the circuit will be underconstrained with specific inputs. Then, we use $\text{AC}^{4}$ to directly solve the systems of quadratic equations with Gröebner basis[25] and check the uniqueness of the output variables.
@@
Figure 3: Decoder(2)
@@
Based on the classification of circuits mentioned above, it is evident that this circuit is quadratic. We begin by treating the variable $inp$ as a coefficient and proceed to solve the undetermined coefficients of each equation. This process yields the solutions [$[\{inp=0\},\{inp=1\}]$. Upon substituting $\{inp=0\}$ into $E$, it results in $E=\{0=0,out_{1}\times(inp-1)=0,success\times(success-1)=0\}$. Notably, it is evident that the $out_{0}$ can take on any value over the finite field, signifying that the circuit is underconstrained.
@@ 4.2 Checking Circuits: Category-Based Algorithms
From above, we divide the overall circuit verification problem space into three types: precisely linear, $k$-coefficient linear, and quadratic separately. In this section, we introduce those algorithms in detail.
Precisely linear. The majority of circuits are precisely linear, so they can be represented by a coefficient matrix $A$ and a constant vector $b$. While it is commonly thought to use the rank of the equation matrix to determine the number of solutions, the solver’s objective is only to verify the uniqueness of output variables. In contrast, rank-reducing methods impose restrictions on the uniqueness of all variables, including temporary variables. Using the Gauss-Jordan method to get the rank of the corresponding matrix is analogous to getting the solution of the system of equations. Consequently, there is no significant difference in the time complexity. The pesudo-code is given in Algorithm 1.
@@
7:         return precisely exactconstrained
@@
Qudratic. A heuristic approach will be used to identify input variables that may cause the circuit to become underconstrained (shown in Algorithm2), aiming to improve accuracy and efficiency. The coefficients of the unknown variables are treated as equations and their solutions will be stored in a frequency table for counting the occurrences of each solution. Then, the solver sorts the solutions based on their frequencies and iteratively substitutes them into the original system of equations to obtain a new system, which is then solved to check if the circuit is underconstrained. If the input variable causing the circuit to become underconstrained cannot be identified, the system of equations will be directly solved with Gröebner basis to verify the constrained situation of the circuit. The pseudo-code of checking quadratic circuits is given in Algorithm3.
Algorithm 2 check circuits by the coefficients of the output variables
procedure GetUndeterminedCoeffSols($K,U,E$)
@@
     for all $eq\in E$ do
         $sols\leftarrow\textsc{SolveUndeterminedCoefficientEq}(E,K,U)$
         if $sols\neq\emptyset$ then
              $UCSolCountTable[sol]++$
@@
Algorithm 3 Algorithm for checking quadratic circuits
procedure CheckQuadraticCircuit($\langle C(\mathcal{X}),E\rangle$)
@@
     else if CheckUniqueness(sols, $O$ then
         return algebraically exactconstrained
@@
$K$-coefficient linear. Although quadratic in nature, these equations are considered linear due to their quadratic terms containing one or more known variables, Hence, they are treated as linear circuits and represented in matrix form to enhance verification efficiency. For these circuits, we still use the previous heuristic approach given in Algorithm 2, to determine if there are special inputs that can cause the underconstrained condition, aiming to enhance efficiency. Afterwards, a $k$-coefficient linear circuit can be succinctly represented by a coefficient matrix $A$ and a constant vector $b$, and checked as a linear case. The pseudo-code of checking $k$-coefficient circuits is given in Algorithm4.
@@
16:         return algebraically exactconstrained
@@ 4.3 Complexity Analysis
This section discusses the time complexity of each category. First, for linear or part of $K$-coefficient linear circuits, the Gauss-Jordan solution yields a complexity of $O(n^{3})$ [2]. As we discovered that the majority of cases are precisely linear circuits based on a statistical analysis of the Circomlib [23], We expect our algebraic-based method to introduce a large improvement in performance compared with SMT solvers, as SMT solvers tend to convert constraints into Conjunctive Normal Form (CNF) which ignores potential mathematical transformation and simplification within equation systems. When considering a quadratic equations system over a finite domain, the time complexity is proven to be NP-hard [17]. Even with complex cases, quadratic solvers offer a practical approach.
4.4 Optimization and Implementation
Pre-implementation. Before solving the equations, it is crucial to first examine whether there are any output variables that are not present in any constraint, as the presence of such a variable indicates that the circuit is underconstrained. Once identified, these variables can then be addressed appropriately. Furthermore, it is important to note the existence of equations with no unknown terms. These equations should be filtered out as they do not contribute to the determination of the output variables’ uniqueness. In addition, it is necessary to carefully assess the type of each equation to establish the type of the circuit being analyzed ultimately. Moreover, in scenarios involving $k$-coefficients, it is possible for the result to be algebraically overconstrained, thereby requiring a methodical approach to narrow down its range. To achieve this, the first step involves checking for linear input variables and treating them as unknown in order to attain more accurate solutions. This systematic process enhances the precision and reliability of the solutions obtained.
Binary Optimization. We improved the efficiency of checking the binary circuits by first encoding them into binary format over a finite field. Then, we implemented a detector to verify whether a quadratic circuit is a binary circuit by ensuring that all unknown variables adhere to the form $u\times(u-1)=0$. This binary checking process was employed to validate the circuits further.
Handling of finite field operations. In our experiments, we observed that only a small number of binary circuits would experience overflow in the finite domain. Despite the availability of relatively efficient algorithms for polynomial operations over finite fields, such as Berlekamp’s algorithm[10], the large characteristic of the finite field of the circom can result in less efficient operations. As a result, we opted to employ integers in the practical implementation. Additionally, we utilized a table to strategically map overflow values and their corresponding original values within a select few binary circuit programs. This approach was implemented to enhance the overall efficiency of the system.
@@ 5 Evaluations
5.1 The Conditions of the Data Set
The size condition Table 1 presents the size condition of the benchmark set. Notably, within the benchmark set Core, there are three circuits with over 200,000 constraints, which are believed to be unsolvable using standard computers. Interestingly, in these three cases, only Sha256compression results in an out-of-memory exception due to the presence of around 20,000 unknown variables, resulting in the creation of a very large matrix. In contrast, the other two cases each have only one known variable, making the circuits easier to check.
Table 1: The size condition of the benchmark sets
@@ 5.1 The Conditions of the Data Set
We categorize circuits based on their constraint numbers in Table 2, considering those with a constraint number less than 100 as ”small” circuits, those with a constraint number in the range $[100,1000)$ as ”medium” circuits, and those with a constraint number in the range $[1000,\infty)$ as ”large” circuits. This classification, which may be rudimentary, does not consider output signals. It is evident that small circuits can be solved directly, whereas other circuits may require specific restrictions to be solved within an acceptable time frame.
@@
The type condition As shown in the table 4, The benchmark set comprises 176 circuits, distributed as 71.6% precisely linear, 6.2% $K$-coefficient linear, and 22.2% quadratic, mirroring the range of circuit types encountered in real-world applications. Upon further examination of the $K$-coefficient and quadratic cases, it becomes evident that a significant portion of them are binary circuits associated with cryptographic algorithms. Instead of exhaustive arithmetic solutions, focusing on binary optimization strategies offers promising efficiency gains in checking these circuits. For instance, the circuit Num2BitsNeg@bitify_256 is a quadratic circuit designed to convert an integer over a finite field into a 256 bit-vector, with all its unknown variables constrained by $o\times(o-1)=0$, indicating that they are strictly bits with their value limited to 0 or 1. Appropriately identifying these circuits and applying binary methods for solving them represents a more efficient approach compared to arithmetic methods.
5.2 Benchmark Metrics
Checked Ratio. The solved benchmarks are initially categorized into three types: precisely solved, algebraically solved, and unknown. The precisely solved cases are further categorized into three types: precisely exactconstrained, which means that the circuit is safe; precisely overconstrained, indicating that the circuit is overconstrained; and precisely underconstrained, which suggests that the circuit is underconstrained. We omit this metric as $PS$ (Precisely Solved).
On the other hand, we propose a more relaxed metric, $AS$ (Algebraic Solved), by remapping cases that failed to obtain precise results into two new categories: algebraic exactconstrained and algebraic overconstrained as introduced in Section 3.2, to reject overconstrained and underconstrained respectively. Note that $\text{AC}^{4}$ still returns unknown for extremely computationally exhausting cases.
Checking time. In this metric, our emphasis lies on the benchmark data on the parts that both $\text{AC}^{4}$ and Picus can solve, and then proceed to compare the checking time of the two tools.
5.3 Benchmark Results
Checking Time The checking time of the benchmark set is presented in Table 3, with a focus on the relationship between benchmark size and solving difficulty. It is observed that the larger benchmarks present greater difficulty in solving, resulting in longer solution times. However, in some large circuits, despite a high number of constraints, the number of output signals is limited to 1, allowing for a relatively easy check once the file read operation concludes. Nonetheless, there exist extremely large circuits that are characterized by a linear type, yet their extensive constraint size renders the Gaussian-Jordan method challenging to execute, leading to out-of-memory issues. Conversely, smaller cases are relatively straightforward to solve directly but due to limitations, the precise solution is not attainable, resulting in algebraic outcomes.
@@ 5.3 Benchmark Results
5.4 Comparison with Baseline
@@
The comparison of the two tools based on the solved ratio can be found in Fig. 4b. Our advantage lies in precisely linear cases as well as some heuristic methods for $k-$coefficient and binary cases.
Fig. 4b shows the solving time comparison of the two tools. Our advantage lies in solving problems arithmetically rather than encoding them into SMT formulas, making our tools much faster.
5.5 Analysis
Table 3 illustrates a significant correlation between the size of the circuit and the time required for checking. Despite their size, large circuits keep checking times low due to two critical factors. Firstly, the predominance of linear circuits, efficiently addressed by the Gauss-Jordan solver, contributes to the efficiency in solving a majority of cases. Additionally, optimizations targeting binary cases within k-coefficient and quadratic circuits serve to further reduce the solving time despite their slightly larger size.
@@ 6 Related Work
Verification on ZK programs. Circomspect [12] serves as a static analyzer and linter designed for the Circom programming language. Picus [37] employs a static analysis method combined with an SMT solver. Circomspect and Picus share certain similarities with our work. However, Circomspect can only annotate a portion of underconstrained variables. Picus is immune to this issue but remains a low efficiency due to the extensive use of the SMT solver.
@@
Algebra computation in circuit verification. Algebraic computation is a long-standing but niche endeavor in verification. Watanabe [46] utilize Gröbner Basis [19, 18] and polynomial reduction techniques to verify arithmetic circuits in computing and signal processing systems. Lv et al. [32, 31] used a computer-algebra-based approach for the formal verification of hardware. Scholl et al. [40] combined Computer Algebra with Boolean SAT solver to verify divider circuits.
However, we emphasize that most of the work in this field actively reduces program or hardware problems to algebra. To the best of our knowledge, we are the inaugural contributors to the application of algebraic methods in ZKPs.

### v4

@@ 4.1 AC4 Framework
For higher-order circuits, we solve the coefficients and verify if the circuit will be underconstrained with specific inputs. Then, we use $\text{AC}^{4}$ to directly solve the systems of polynomial equations with Gröebner basis[18] and check the uniqueness of the output variables.
@@
Figure 2: Binary to One-Hot Decoder Circuit
@@
Based on the classification of circuits mentioned above, it is evident that this circuit is not linear. To analyze this, we begin by treating the variable $inp$ as a coefficient and proceed to solve for the undetermined coefficients of each equation. This process yields the solutions [$[\{inp=0\},\{inp=1\}]$. When substituting $\{inp=0\}$ into $E$, the resulting equations are $E=\{0=0,out_{1}\times(inp-1)=0,success\times(success-1)=0\}$. Notably, $out_{0}$ can take on any value over the finite field, indicating that the circuit is underconstrained.
4.2 Adaptive Checking with Relaxation and Partial Results
AC4 is designed to provide precise results when analyzing circuits. However, when dealing with large or complex circuits, it might produce an “unknown” outcome due to the scalability. To handle these situations, AC4 relaxes its strict checking conditions and instead returns partial, but still meaningful, results.
[…]
While its primary objective is to provide precise results during circuit analysis, AC4 recognizes the challenges posed by large circuits with numerous parameters. In such scenarios, where a precise result may be unattainable, AC4 adjusts its approach by relaxing the strictness of its checking conditions. This adaptability enables AC4 to yield a meaningful partial result, ensuring that even with complex circuits, users can still gain valuable insights into the circuit’s behavior.
To provide more detailed feedback on unsolvable circuits, AC4 refines its result classification. It introduces two intermediate categories, algebraic exact-constrained and algebraic overconstrained, which offer additional insights. These categories help analysts understand whether a circuit may be underconstrained (too flexible) or overconstrained (no valid solutions) by examining the algebraic structure of the equation matrix.
The relationship between AC4’s results and the actual circuit behavior is also cross-referenced with ground truth. Three categories—precisely underconstrained, precisely exact-constrained, and precisely overconstrained-directly correspond to whether a circuit is under, exact or overconstrained. However, the algebraic categories offer more nuanced information. For example, an algebraically exact-constrained circuit might be either underconstrained or exact-constrained in reality, while algebraic overconstrained cases could actually be exact-constrained or overconstrained.
Unknown results may stem from circuit size, computational complexity, or the inherent NP-hardness of solving polynomial equations over finite fields.
By relaxing the strict checking conditions, AC4 can filter out cases that are definitely under or overconstrained, providing more efficient and useful results even when a full solution isn’t possible.
Figure 3: The category of results from AC4
@@ 4.3 Checking Circuits: Category-Based Algorithms
From above, we divide the overall circuit verification problem space into three types: precisely linear, $k$-coefficient linear, and higher-order separately. In this section, we introduce those algorithms in detail.
Precisely linear. The majority of circuits are precisely linear, so they can be represented by a coefficient matrix $A$ and a constant vector $b$. While it is commonly thought to use the rank of the equation matrix to determine the number of solutions, the solver’s objective is only to verify the uniqueness of output variables. In contrast, rank-reducing methods impose restrictions on the uniqueness of all variables, including temporary variables. Using the Gauss-Jordan method to get the rank of the corresponding matrix is analogous to getting the solution of the system of equations. Consequently, there is no significant difference in the time complexity. The pesudo-code is given in Algorithm
@@
7:         return precisely exact-constrained
@@
Higher-Order. A heuristic approach will be used to identify input variables that may cause the circuit to become underconstrained (shown in Algorithm3), aiming to improve accuracy and efficiency. The coefficients of the unknown variables are treated as equations and their solutions will be stored in a frequency table for counting the occurrences of each solution. Then, the solver sorts the solutions based on their frequencies and iteratively substitutes them into the original system of equations to obtain a new system, which is then solved to check if the circuit is underconstrained. If the input variable causing the circuit to become underconstrained cannot be identified, the system of equations will be directly solved with Gröbner basis to verify the constrained situation of the circuit. The pseudo-code of checking polynomial circuits is given in Algorithm2.
Algorithm 2 Check multiple solutions by analyzing coefficients of parameters
procedure CheckMultiSolutionsByCoeffs($K,U,E$)
@@
     for all $eq\in E$ do $\triangleright$ iterate all the equations
         $sols\leftarrow\textsc{SolveUndeterminedCoefficientEq}(eq,K,U)$
         if $sols\neq\emptyset$ then $\triangleright$ If there exists solution
              for all $sol\in sols$ do
                  $UCSolCountTable[sol]\leftarrow UCSolCountTable[sol]+1$
              end for
@@
Algorithm 3 Algorithm for checking higher-order circuits
procedure CheckHigherCircuit($\langle C(\mathcal{X}),E\rangle$)
@@
     else if CheckUniqueness(sols, $O$) then
         return algebraically exact-constrained
@@
Although in nature, these equations are considered linear due to their higher-order terms containing one or more known variables, Hence, they are treated as linear circuits and represented in matrix form to enhance verification efficiency. For these circuits, we still use the previous heuristic approach given in Algorithm 2, to determine if there are special inputs that can cause the underconstrained condition, aiming to enhance efficiency. Afterwards, a $k$-coefficient linear circuit can be succinctly represented by a coefficient matrix $A$ and a constant vector $b$, and checked as a linear case. The pseudo-code of checking $k$-coefficient circuits is given in Algorithm4.
@@
16:         return algebraically exact-constrained
@@ 4.4 Complexity Analysis
This section examines the time complexity associated with each category of circuits. For linear circuits, including those characterized by $K$-coefficient linear circuits, the Gauss-Jordan elimination method has a time complexity of $O(n^{3})$ [3]. Our statistical analysis of Circomlib [32] indicates that the majority of cases involve linear circuits. Consequently, we anticipate that our algebraic-based method will significantly enhance performance compared to SMT solvers. In the case of higher-order circuits, we employed the $F_{5}$ algorithm, which is designed for computing Gröbner bases.[23] This algorithm typically exhibits $O(2^{n})$ or $O(2^{2^{n}})$ time complexity in the worst-case scenarios. However, the $F_{5}$ algorithm demonstrates substantial improvements over its predecessors, such as Buchberger’s algorithm[18] and the $F_{4}$ algorithm.
@@ 5.1 Environment Setting and Metrics
Furthermore, we established a benchmark to evaluate the effectiveness and accuracy of AC4 and the halo2-analyzer[31], as implemented in the study by Soureshjani et al. (2023) [52]. This benchmark was constructed using a dataset of classical halo2 circuits, which were collected from various GitHub projects.
The primary benchmark focuses on circom datasets, while our results also demonstrate improvements on halo2 datasets compared to other tools.
Benchmark Metrics Checked Ratio. The solved benchmarks are initially categorized into three types: precisely solved, algebraically solved, and unknown. The precisely solved cases are further categorized into three types: precisely exact-constrained, which means that the circuit is safe; precisely overconstrained, indicating that the circuit is overconstrained; and precisely underconstrained, which suggests that the circuit is underconstrained. We omit this metric as $PS$ (Precisely Solved).
On the other hand, we propose a more relaxed metric, $AS$ (Algebraic Solved), by remapping cases that failed to obtain precise results into two new categories: algebraic exact-constrained and algebraic overconstrained, to reject overconstrained and underconstrained respectively. Note that $\text{AC}^{4}$ still returns unknown for extremely computationally exhausting cases.
Checking time. In this metric, our emphasis lies on the benchmark data on the parts that both $\text{AC}^{4}$ and other tools can solve, and then proceed to compare the checking time of the tools.
5.2 Detailed Evaluation over Circom Circuits
The Conditions of the Data Set over Circom
The size condition. Table 1 presents the size condition of the benchmark set. Notably, within the benchmark set Core, there are three circuits with over 200,000 constraints, which are believed to be unsolvable using standard computers. Interestingly, in these three cases, only Sha256compression results in an out-of-memory exception due to the presence of around 20,000 unknown variables, resulting in the creation of a very large matrix. In contrast, the other two cases each have only one known variable, making the circuits easier to check.
Table 1: The size condition of the Circom benchmark sets
@@ 5.2 Detailed Evaluation over Circom Circuits
We categorize circuits based on their constraint numbers in Table 2, considering those with a constraint number less than 100 as “small” circuits, those with a constraint number in the range $[100,1000)$ as “medium” circuits, and those with a constraint number in the range $[1000,\infty)$ as “large” circuits. This classification, which may be rudimentary, does not consider output signals. It is evident that small circuits can be solved directly, whereas other circuits may require specific restrictions to be solved within an acceptable time frame.
@@
The type condition. As shown in the table 4, The benchmark set comprises 176 circuits, distributed as 71.6% precisely linear, 6.2% $K$-coefficient linear, and 22.2% quadratic, mirroring the range of circuit types encountered in real-world applications. Upon further examination of the $K$-coefficient and quadratic cases, it becomes evident that a significant portion of them are binary circuits associated with cryptographic algorithms. Instead of exhaustive arithmetic solutions, focusing on binary optimization strategies offers promising efficiency gains in checking these circuits. For instance, the circuit Num2BitsNeg@bitify_256 is a quadratic circuit designed to convert an integer over a finite field into a 256 bit-vector, with all its unknown variables constrained by $o\times(o-1)=0$, indicating that they are strictly bits with their value limited to 0 or 1. Appropriately identifying these circuits and applying binary methods for solving them represents a more efficient approach compared to arithmetic methods.
Benchmark Results over Circom
Checking time. The checking time of the benchmark set is presented in Table 3, with a focus on the relationship between benchmark size and solving difficulty. It is observed that the larger benchmarks present greater difficulty in solving, resulting in longer solution times. However, in some large circuits, despite a high number of constraints, the number of output signals is limited to 1, allowing for a relatively easy check once the file read operation concludes. Nonetheless, there exist extremely large circuits that are characterized by a linear type, yet their extensive constraint size renders the Gaussian-Jordan method challenging to execute, leading to out-of-memory issues. Conversely, smaller cases are relatively straightforward to solve directly but due to limitations, the precise solution is not attainable, resulting in algebraic outcomes.
@@
Comparison with Baseline
@@
The comparison of the two tools based on the solved ratio can be found in Fig. 4b. Our advantage lies in precisely linear cases as well as some heuristic methods for $k-$coefficient and binary cases.Fig. 4b shows the solving time comparison of the two tools.
5.3 Evaluation over Halo2 Circuits
Given the current dearth of comprehensive research on halo2 circuits and the absence of template circuits provided by the official Zcash organization, we undertook the task of extracting 10 of the most frequently utilized halo2 circuit types from GitHub repositories dedicated to halo2 projects. These extractions were subsequently employed to construct a relevant dataset. Utilizing this custom-curated dataset, we conducted benchmark analyses in conjunction with the halo2-analyzer.
As shown in figure 5, AC4 demonstrates a enhancement in efficiency and accuracy compared to halo2-analyzer, exhibiting a 26.73% improvement in the verification time and 10% improvement in the checked ratio for halo2 circuits.
(a) Comparison between the checking time of $\text{AC}^{4}$ and halo2-analyzer.
(b) Comparison between checked ratio of $\text{AC}^{4}$ and Halo2-analyzer
Figure 5: The checking time and checked ratio comparisons between $\text{AC}^{4}$ and halo2-analyzer
5.4 Evaluation Result Analysis
Table 3 illustrates a significant correlation between the size of the circuit and the time required for checking. Despite their size, large circuits keep checking times low due to two critical factors. Firstly, the predominance of linear circuits, efficiently addressed by the Gauss-Jordan solver, contributes to the efficiency in solving a majority of cases. Additionally, optimizations targeting binary cases within k-coefficient and higher-order circuits serve to further reduce the solving time despite their slightly larger size.
@@ 5.4 Evaluation Result Analysis
In the context of quadratic cases, Picus and halo2-analyzer employ the cvc5-ff-range solver. This solver first implements finite field theory and then utilizes Buchberger’s algorithm along with triangular decomposition to simplify polynomial equations. To address the problem, it further applies Collins’ cylindrical algebraic decomposition (CAD) [22] [43]. It is important to note that the time complexity of CAD is $O(2^{2^{n}})$, which can be prohibitive in many scenarios involving large quadratic cases. Additionally, the $F_{5}$ algorithm demonstrates superior performance over modular integers in various cryptographic challenges.
@@
Analysis over halo2 circuits Halo2-analyzer employs cvc5-ff to verify circuit uniqueness, giving our tool an advantage in solving large-scale polynomial equation systems.
@@ 6 Related Work
Verification on ZK programs. Circomspect [14] serves as a static analyzer and linter designed for the Circom programming language. Picus [46] employs a static analysis method combined with an SMT solver. Circomspect and Picus share certain similarities with our work. However, Circomspect can only annotate a portion of underconstrained variables.
@@
Algebra computation in circuit verification. Algebraic computation is a long-standing but niche endeavor in verification. Watanabe [56] utilized Gröbner Basis [24, 23] and polynomial reduction techniques to verify arithmetic circuits in computing and signal processing systems. Similarly, Lv et al. [40, 39] employed a computer-algebra-based approach for the formal verification of hardware. Scholl et al. [49] advanced this field by combining Computer Algebra with a Boolean SAT solver to verify divider circuits. In the context of Zero Knowledge Proof (ZKP) compilers, Alex et al. [45] presented a partial verification of a field-blasting compiler pass. Furthermore, Thomas et al. [30] introduced a novel automated reasoning method for determining the satisfiability of non-linear equation systems over finite fields. Additionally, Thomas et al. [29] enhanced the Yices2 SMT solver by implementing a new MCSat-based reasoning engine for finite fields.
Alex et al. [43] employed the C++ CLN library [17] to implement finite fields and utilized COCOALib’s [1] implementation of the Buchberger algorithm for preprocessing equations, subsequently applying CAD algorithm[22] to solve them and presents a new SMT solver for finite field equations that improves efficiency on cryptosystem verification tasks by using multiple simplified Gröbner bases and specialized propagation algorithms[44].
We emphasize that much of the work in this field reduces program or hardware issues to algebra. To our knowledge, we are the first to apply computer algebra methods in ZKPs.